OpenStudy (anonymous):

Find the product and express your answer in scientific notation. (5 x 106)(8.2 x 10-9)

OpenStudy (kkutie7):

\[(5*10^6)(8.2*10^{-9})\]

OpenStudy (kkutie7):

think of it like this: \[(5x^6)(8.2x^{-9})\] what woul you do here?

OpenStudy (kkutie7):

if you don't get it I'll help you just have to let me know

OpenStudy (anonymous):

all I have to do is multiply?

OpenStudy (photon336):

@dezmonique still need help?

OpenStudy (anonymous):

yes I do @Photon336

OpenStudy (photon336):

so this is scientific notation right

OpenStudy (anonymous):

yes

OpenStudy (photon336):

so when we multiply exponents we add them first

OpenStudy (photon336):

\[10^{6}*10^{-9} = 10^{6+-9}\]

OpenStudy (photon336):

so let's ignore the numbers 5 and 8.2 for now

OpenStudy (photon336):

@dezmonique whenever we have two bases and we have to multiply them. we add the exponents. \(\color{blue}{\text{Originally Posted by}}\) @Photon336 \[10^{6}*10^{-9} = 10^{6+-9}\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

so it would be 10^-3 right?

OpenStudy (photon336):

yes

OpenStudy (photon336):

now let's multiply 5 and 8.2

OpenStudy (anonymous):

okay its 41 right?

OpenStudy (photon336):

so now let's put them together \[41*10^{-3} = ? \]

OpenStudy (anonymous):

is it in a decimal form?

OpenStudy (photon336):

so we're almost done

OpenStudy (photon336):

we need to express our last number in prober scientific notation 41*10^{-3} isn't acceptable so here is what we can do \[41 = 4.1*10^{1}*10^{-3}\]

OpenStudy (photon336):

\[4.1*10^{-2}\]

OpenStudy (anonymous):

is that how we would write it in scientific notation?\[4.1*10^-2\]

OpenStudy (photon336):

yep :)

OpenStudy (anonymous):

okay thank you :) that helped i was so confused in how it was suppose to be done

OpenStudy (photon336):

try another problem

OpenStudy (photon336):

\[(2*10^{3})(3*10^{2})\]

OpenStudy (photon336):

@dezmonique

OpenStudy (anonymous):

this is what i got so far, \[60^5\]

OpenStudy (anonymous):

would it be 6*10^6? @Photon336

OpenStudy (photon336):

almost

OpenStudy (photon336):

2*3 = 6 and 2+3 = 5 6*10^5

OpenStudy (anonymous):

i thought it was to the 5th power to, but then i put to the 6th power

OpenStudy (photon336):

okay try this one \[(3*10^{2})(3*10^{6}) = ?\]

OpenStudy (anonymous):

\[9*10^8\]

OpenStudy (photon336):

excellent

OpenStudy (photon336):

one more

OpenStudy (anonymous):

okay

OpenStudy (photon336):

\[(4*10^{3})(3*10^{2})\]

OpenStudy (anonymous):

\[12*10^5\]

OpenStudy (photon336):

yep

OpenStudy (photon336):

or \[1.2*10^{6}\]

OpenStudy (anonymous):

how do you get that answer?

OpenStudy (photon336):

so remember scientific notation is written the number must be less than 10

OpenStudy (photon336):

We can re-write \[12 = 1.2*10^{1}\]

OpenStudy (anonymous):

oh okay i get it

OpenStudy (photon336):

so it becomes this \[(1.2*10^{1})(10^{5})\]

OpenStudy (photon336):

which is just \[(1.2*10^{6})\]

OpenStudy (anonymous):

oh, got it

OpenStudy (photon336):

last one \[(4*10^{2})(3*10^{3})\]

OpenStudy (photon336):

anyway i g2g @dezmonique try that last one on your own

OpenStudy (anonymous):

okay thank you for your help